Daily Report

CPUC/ANF Environmental Monitoring Activities

Pacific Pipeline Construction Project

Aspen Environmental Group Wednesday, November 18, 1998

The following is a brief summary of the environmental monitoring activities that were completed on November 18, 1998. This summary is based on reports directly received from the field and has not been edited. If you have any questions please contact (818) 597-3407. Please note the item number and the correction that should be made. This report is intended to serve as a communication document and as an instrument for quality control.

Construction Highlights

11/18/98 (Wednesday)

Segment 1 report for 11/18/98: On Lebec Road at the Warbler skip, clearing began from EMP's 21.56 to 21.58. A track hoe began excavation at EMP 21.50 working south. On Tejon Bench welding and tie-in occurred from 18.14 to 18.12. X-ray occurred from 16.72 to 16.70. In Grapevine Canyon, trench was excavated to EMP 15.50, pipe bending occurred to 15.93, welding from 16.11 to 16.08. Bell holes were being excavated from 15.76 to 15.70. Coating occurred to 16.32. At Grapevine Pump Station, electrical work continued at the substation and electrical building. Pipe-fitting and excavation for pipe placement also occurred. Tie-in took place on both sides of the I-5/Grapevine bore with pipe being placed on the east side from EMP's 14.68 to 14.67 and welding, x-ray and coating occurring on the west side from EMP's from 14.65 to 14.64. On Hwy. 166, two track hoes continued to excavated, one working east from EMP .61 and the other working west from EMP .74. Electrical work continued at Emidio Pump Station.

Segment 2 (ANF to Sylmar) Report for 11/18/98: 2B (Castaic to Sylmar): Boring at the Santa Clara River directional drill at Feedmill Rd reached 800 feet. 500 feet of boring remain to be completed. A pump system was set up to aid control of the frac-outs. Welding, x-ray, and coating continued north of Sierra Hwy. South of Sierra Hwy, 800 feet of F/O conduit was laid to EMP 72.76, 400 feet of backfill completed, 200 feet of shading to 72.85, and 200 feet of asphalt applied to 73.05.

Segment 3A & 3C (Sylmar to Los Angeles): Report for 18 November 1998: 3A: Fiber optic splicing was done near the valve under the 2 Freeway today. Fiber optic cable was pulled in Glendale, between Highland and Colorado Streets. Backfill crews repaired a large sunken area of the right of way just off of Linden Street, north of the valve. Other backfill crews neared completion of the south side of the Hollywood Way Bridge crossing. Pipe was delivered to the right of way south of Tujunga Wash, while ditch crews dug out the tie in areas. The bore began at Tuxford Street, including a small crew working on the exit hole to the north. Welding, x-ray, and coating continued to work on the bore tie in at Van Nuys Boulevard. The remaining brackets were installed on the Pacoima Wash Bridge crossing. Electrical work was done at the valve just south of Highland Street in Glendale. 3C: Welding, x-ray, and coating crews worked on the tie in area near the aqueduct. The Hubbard bore tie in continued. More slurry was poured as backfill in San Fernando Road, near Sepulveda Boulevard. Crews placed wooden railing and dug out for construction of the aqueduct channel near Balboa Boulevard.

Segment 4B(El Segundo to Lynwood) 11-18-98: Construction activities continued at El Segundo Station inside the Chevron refinery ESEMP 11.00 and the 90 degree tie-in (out side the refinery east side of Sepulveda ESEMP 10.78) has been completed. Backfilling activities continued at Western bore pit ESEMP 5.56. The crew inside Lynwood station were working on the south wall. Paving crew was working at 109 Th St. ESEMP 1.23. The contractor had problem whit the backfill at Northrop today. Apparently were they had backfill several months ago, has sink-in more than a feet, so they started the backfilling process again.

Segment 5 (Lynwood to Long Beach) Report for 18 November 1998: Directional drill under Highway 91/Compton Creek is progressing slowly through clay -- less than 30 feet per hour. Two more joints of pipe were pushed through the casing at Wilmington. Crews have begun excavation for the bore under Alameda Street to tie-in to Wilmington bore pipe.

Inspection Summary

In the past 24 hours, 5 inspection event(s) were reported for 4 spread(s). Of these inspections, 0 were non-compliance event(s). Also 0 non-compliance resolution report(s) was resolved.

Non-Compliance Reports (NCR's) Issued

No new non-compliance reports were issued during the last reporting period.

Non-Compliance Resolution Reports (NCRR's) Issued

No new non-compliance reports were issued during the last reporting period.

Summary of Follow-up Items for Today

No follow-up items were reported for today.

Back to Daily Reports

Back to MMCRP Home Page